### v2

@@ A.5 Agglomerative clustering with 10% noise
Appendix B Experiments
@@
B.1 KNN experiments
B.1.1 Without noise
| Euclidean | 2 | 0.89 | 0.89 | 2 | 1.00 | 1.00 | 2 | 0.68 | 0.65 | 1 | 0.56 | 0.57 | 10 | 0.84 | 0.88 | 8 | 0.75 | 0.75 | 9 | 0.91 | 0.90 | 2 | 0.60 | 0.37 |
| Manhattan | 1 | 0.91 | 0.90 | 1 | 1.00 | 1.00 | 2 | 0.70 | 0.66 | 4 | 0.56 | 0.58 | 8 | 0.86 | 0.87 | 10 | 0.74 | 0.76 | 9 | 0.91 | 0.91 | 2 | 0.60 | 0.37 |
| Minkowski | 2 | 0.88 | 0.87 | 2 | 1.00 | 1.00 | 2 | 0.68 | 0.65 | 3 | 0.56 | 0.50 | 10 | 0.85 | 0.88 | 8 | 0.76 | 0.76 | 9 | 0.90 | 0.89 | 2 | 0.59 | 0.36 |
| Cosine | 2 | 0.89 | 0.89 | 1 | 0.99 | 0.99 | 8 | 0.69 | 0.69 | 6 | 0.52 | 0.55 | 3 | 0.93 | 0.90 | 2 | 0.74 | 0.76 | 10 | 0.88 | 0.88 | 1 | 0.55 | 0.36 |
| Hassanat | 2 | 0.92 | 0.92 | 1 | 1.00 | 1.00 | 10 | 0.73 | 0.71 | 3 | 0.67 | 0.67 | 3 | 0.84 | 0.82 | 4 | 0.74 | 0.76 | 10 | 0.91 | 0.90 | 2 | 0.58 | 0.37 |
| Canberra | 6 | 0.92 | 0.91 | 1 | 0.93 | 0.93 | 3 | 0.70 | 0.70 | 9 | 0.66 | 0.71 | 3 | 0.84 | 0.79 | 5 | 0.73 | 0.75 | 6 | 0.92 | 0.90 | 2 | 0.58 | 0.36 |
| Hellinger | 2 | 0.90 | 0.90 | 1 | 0.95 | 0.95 | 8 | 0.70 | 0.69 | 8 | 0.65 | 0.69 | 8 | 0.87 | 0.88 | 7 | 0.73 | 0.76 | 5 | 0.93 | 0.93 | 2 | 0.60 | 0.36 |
| Jensen-Shannon | 2 | 0.78 | 0.40 | 1 | 0.89 | 0.89 | 10 | 0.71 | 0.70 | 10 | 0.63 | 0.67 | 1 | 0.94 | 0.94 | 7 | 0.74 | 0.75 | 9 | 0.89 | 0.88 | 2 | 0.72 | 0.06 |
| PearsonChi2 | 1 | 0.41 | 0.64 | 1 | 0.95 | 0.95 | 8 | 0.73 | 0.71 | 7 | 0.63 | 0.67 | 8 | 0.87 | 0.89 | 8 | 0.72 | 0.75 | 7 | 0.92 | 0.92 | 2 | 0.58 | 0.36 |
| Vicis Symmetric | 7 | 0.81 | 0.76 | 1 | 0.92 | 0.92 | 8 | 0.70 | 0.70 | 10 | 0.62 | 0.67 | 8 | 0.85 | 0.88 | 5 | 0.74 | 0.71 | 8 | 0.91 | 0.91 | 2 | 0.60 | 0.35 |
| Gini | 2 2 | 0.85 | 0.84 | 1 2 | 1.00 | 1.00 | 6 2 | 0.71 | 0.70 | 1 2 | 0.61 | 0.61 | 10 2 | 0.85 | 0.89 | 5 2 | 0.71 | 0.69 | 5 | 0.92 | 0.92 | 2 | 0.58 | 0.37 |
| Generalized Gini | 2 4.89 | 0.90 | 0.89 | 1 1.96 | 1.00 | 1.00 | 7 3.35 | 0.73 | 0.73 | 1 2.12 | 0.63 | 0.63 | 7 5.09 | 0.87 | 0.89 | 5 5.09 | 0.72 | 0.75 | 8 5.09 | 0.93 | 0.93 | 2 2.67 | 0.58 | 0.37 |
B.1.2 With noise at level 5%
@@
B.1.3 With noise at level 10%
@@
B.2 KMeans experiments
B.2.1 Without noise
| Gini Generalized | 1.18 | 0.4915 | 0.4161 | 4.88 | 0.6339 | 0.6363 | 2.48 | 0.6252 | 0.5650 | 1.72 | 0.6186 | 0.3495 | 1.84 | 0.8214 | 0.5520 | 2.14 | 0.4772 | 0.3409 | 2.36 | 0.4466 | 0.1273 | 1.62 | 0.7052 | 0.0759 |
| Gini | - | 0.3897 | 0.3535 | - | 0.5822 | 0.5816 | - | 0.4645 | 0.5360 | - | 0.4916 | 0.3330 | - | 0.7131 | 0.5408 | - | 0.4082 | 0.3475 | - | 0.1467 | 0.1273 | - | 0.6828 | 0.1174 |
| Euclidean | - | 0.3147 | 0.2882 | - | 0.6077 | 0.6115 | - | 0.4123 | 0.5130 | - | 0.4239 | 0.2672 | - | 0.6977 | 0.5296 | - | 0.3183 | 0.3213 | - | 0.5445 | 0.1227 | - | 0.4288 | 0.0155 |
| Manhattan | - | 0.3586 | 0.3280 | - | 0.6083 | 0.6123 | - | 0.4123 | 0.5130 | - | 0.4113 | 0.2573 | - | 0.7204 | 0.5728 | - | 0.3247 | 0.3443 | - | 0.5992 | 0.0977 | - | 0.4288 | 0.0155 |
| Minkowski | 3.0 | 0.3200 | 0.2910 | 3.0 | 0.6085 | 0.6123 | 3.0 | 0.4123 | 0.5130 | 3.0 | 0.4254 | 0.2605 | 3.0 | 0.6475 | 0.4928 | 3.0 | 0.3208 | 0.2909 | 3.0 | 0.5315 | 0.1386 | 3.0 | 0.4288 | 0.0155 |
| Cosine | - | 0.4013 | 0.3621 | - | 0.5974 | 0.6006 | - | 0.3084 | 0.4840 | - | 0.5479 | 0.2934 | - | 0.7292 | 0.5088 | - | 0.3259 | 0.3770 | - | 0.5952 | 0.4886 | - | 0.6376 | 0.1575 |
| Canberra | - | 0.5419 | 0.4505 | - | 0.5901 | 0.5948 | - | 0.5435 | 0.5250 | - | 0.5651 | 0.2013 | - | 0.4913 | 0.3744 | - | 0.6109 | 0.4388 | - | 0.2348 | 0.1136 | - | 0.6864 | 0.1798 |
| Hellinger | - | 0.5754 | 0.5021 | - | 0.6015 | 0.5868 | - | 0.3364 | 0.4740 | - | 0.4159 | 0.1985 | - | 0.5745 | 0.3760 | - | 0.4612 | 0.4304 | - | 0.3319 | 0.2273 | - | 0.7150 | 0.0414 |
| Jensen-Shannon | - | 0.1289 | 0.3590 | - | 0.3085 | 0.5554 | - | 0.3095 | 0.5050 | - | 0.2548 | 0.0363 | - | 0.4687 | 0.1808 | - | 0.2361 | 0.0230 | - | 0.2621 | 0.1523 | - | 0.7150 | 0.0588 |
| PearsonChi2 | - | 0.1289 | 0.3590 | - | 0.5043 | 0.5094 | - | 0.2789 | 0.4960 | - | 0.4802 | 0.2644 | - | 0.4578 | 0.3616 | - | 0.5301 | 0.4550 | - | 0.5397 | 0.4955 | - | 0.6623 | 0.1678 |
| VicisSymmetric1 | - | 0.5412 | 0.5070 | - | 0.5225 | 0.5109 | - | 0.2954 | 0.4970 | - | 0.3202 | 0.1551 | - | 0.4240 | 0.3840 | - | 0.5324 | 0.4189 | - | 0.3484 | 0.1250 | - | 0.5064 | 0.4039 |
| Hassanat | - | 0.4226 | 0.3821 | - | 0.5556 | 0.5517 | - | 0.5754 | 0.5750 | - | 0.5199 | 0.4033 | - | 0.5938 | 0.4464 | - | 0.5128 | 0.4769 | - | 0.3946 | 0.3250 | - | 0.6958 | 0.1400 |

### v3

@@ 5.3 Illustration on Agglomerative Hierarchical Clustering
In this experiment, Ward’s method has been added. It minimizes the within-cluster variance at each step of the clustering process. To be precise, it aims to minimize the increase in the sum of squared within-cluster distances (within-cluster inertia) when two clusters are merged. As can be seen in Table 13, Ward’s technique achieves the highest recall, ranking in the top position 7 times with an average rank of 2.62. The generalized Gini prametric ranks second (see also Tables 27-30 in Supplementary Materials [27]).
6 Conclusion
The generalized Gini prametric, while not being a true distance, satisfies two invariance properties, making it robust to outliers. In particular, it is linearly invariant (similar to the Euclidean distance) and rank invariant. Accordingly, monotonic increasing transformations preserve ranks, ensuring that the Gini prametric exhibits small variations. Consequently, when the data contain outliers or measurement errors, the generalized Gini prametric helps achieve robust results.
The experiments on the UCI datasets highlight the limitations of the generalized Gini prametric in unsupervised tasks. In particular, a strategy for selecting the hyper-parameter, such as using the silhouette score, must be chosen without prior knowledge of the quality of the results. This remains an open problem.
Supplementary Materials
Proof of Proposition 4
The last inequality comes from the algorithm itself. Indeed, if , this is because at iteration we have . Now, we must prove that taking the sum over on the last expression implies:
with . Taking the derivative of with respect to , and imposing that for small variations of ranks remain constant, yields
Using the scalar product this is equivalent to:
Thus,
This implies,
Experiments
@@ 6 Conclusion
KNN experiments
Without noise
Ionosphere Banknote German Heart Balance Haberman Wholesale Indian Liver Patient Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Euclidean 2 0.89 0.89 2 1.00 1.00 2 0.68 0.65 1 0.56 0.57 10 0.84 0.88 8 0.75 0.75 9 0.91 0.90 2 0.60 0.37 Manhattan 1 0.91 0.90 1 1.00 1.00 2 0.70 0.66 4 0.56 0.58 8 0.86 0.87 10 0.74 0.76 9 0.91 0.91 2 0.60 0.37 Minkowski 2 0.88 0.87 2 1.00 1.00 2 0.68 0.65 3 0.56 0.50 10 0.85 0.88 8 0.76 0.76 9 0.90 0.89 2 0.59 0.36 Cosine 2 0.89 0.89 1 0.99 0.99 8 0.69 0.69 6 0.52 0.55 3 0.93 0.90 2 0.74 0.76 10 0.88 0.88 1 0.55 0.36 Hassanat 2 0.92 0.92 1 1.00 1.00 10 0.73 0.71 3 0.67 0.67 3 0.84 0.82 4 0.74 0.76 10 0.91 0.90 2 0.58 0.37 Canberra 6 0.92 0.91 1 0.93 0.93 3 0.70 0.70 9 0.66 0.71 3 0.84 0.79 5 0.73 0.75 6 0.92 0.90 2 0.58 0.36 Hellinger 2 0.90 0.90 1 0.95 0.95 8 0.70 0.69 8 0.65 0.69 8 0.87 0.88 7 0.73 0.76 5 0.93 0.93 2 0.60 0.36 Jensen-Shannon 2 0.78 0.40 1 0.89 0.89 10 0.71 0.70 10 0.63 0.67 1 0.94 0.94 7 0.74 0.75 9 0.89 0.88 2 0.72 0.06 PearsonChi2 1 0.41 0.64 1 0.95 0.95 8 0.73 0.71 7 0.63 0.67 8 0.87 0.89 8 0.72 0.75 7 0.92 0.92 2 0.58 0.36 VicisSymmetric1 7 0.81 0.76 1 0.92 0.92 8 0.70 0.70 10 0.62 0.67 8 0.85 0.88 5 0.74 0.71 8 0.91 0.91 2 0.60 0.35 Gini 2 2 0.85 0.84 1 2 1.00 1.00 6 2 0.71 0.70 1 2 0.61 0.61 10 2 0.85 0.89 5 2 0.71 0.69 5 0.92 0.92 2 0.58 0.37 Generalized Gini 2 4.89 0.90 0.89 1 1.96 1.00 1.00 7 3.35 0.73 0.73 1 2.12 0.63 0.63 7 5.09 0.87 0.89 5 5.09 0.72 0.75 8 5.09 0.93 0.93 2 2.67 0.58 0.37
Iris Wine Breast-Cancer Sonar QSAR Vehicle Australian Glass Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Euclidean 5 - 0.97 0.97 1 - 0.71 0.71 10 - 0.93 0.93 1 - 0.84 0.83 6 - 0.64 0.65 1 - 0.87 0.87 6 - 0.70 0.69 3 - 0.74 0.72 Manhattan 1 - 0.96 0.96 1 - 0.77 0.77 6 - 0.94 0.94 1 - 0.86 0.86 9 - 0.64 0.65 1 - 0.90 0.90 8 - 0.75 0.73 1 - 0.75 0.71 Minkowski 5 - 0.97 0.97 1 - 0.68 0.69 10 - 0.93 0.93 3 - 0.84 0.83 8 - 0.63 0.64 1 - 0.85 0.85 10 - 0.70 0.67 2 - 0.68 0.65 Cosine 7 - 0.98 0.97 1 - 0.84 0.84 7 - 0.93 0.92 1 - 0.86 0.85 9 - 0.63 0.63 1 - 0.89 0.88 9 - 0.72 0.72 8 - 0.69 0.69 Hassanat 10 - 0.95 0.95 1 - 0.96 0.95 7 - 0.96 0.96 1 - 0.87 0.87 8 - 0.63 0.64 7 - 0.89 0.89 8 - 0.86 0.84 4 - 0.74 0.68 Canberra 9 - 0.95 0.95 1 - 0.97 0.97 6 - 0.96 0.95 1 - 0.85 0.83 7 - 0.65 0.65 1 - 0.85 0.85 6 - 0.87 0.86 2 - 0.67 0.68 Hellinger 2 - 0.97 0.97 4 - 0.91 0.92 4 - 0.95 0.95 1 - 0.87 0.86 6 - 0.64 0.65 1 - 0.86 0.86 2 - 0.77 0.76 1 - 0.69 0.69 Gini 4 2 0.97 0.97 5 2 0.90 0.89 5 2 0.95 0.95 1 2 0.87 0.87 7 - 0.66 0.67 1 2 0.93 0.93 6 2 0.77 0.75 3 2 0.73 0.71 Generalized Gini 7 1.1 0.98 0.98 3 4.45 0.91 0.91 4 3.11 0.95 0.95 1 2.6 0.88 0.88 7 2.39 0.67 0.67 1 3.48 0.93 0.93 5 3.45 0.79 0.79 3 1.37 0.77 0.74 JensenShannon 3 - 0.98 0.96 10 - 0.95 0.94 3 - 0.93 0.93 1 - 0.87 0.86 9 - 0.64 0.65 9 - 0.86 0.86 9 - 0.77 0.78 1 - 0.67 0.66 PearsonChi2 2 - 0.97 0.97 4 - 0.92 0.92 4 - 0.95 0.95 1 - 0.86 0.86 2 - 0.57 0.59 7 - 0.76 0.75 7 - 0.76 0.75 1 - 0.68 0.67 VicisSymmetric1 5 - 0.97 0.95 1 - 0.98 0.98 5 - 0.96 0.96 1 - 0.79 0.78 9 - 0.65 0.65 5 - 0.78 0.78 8 - 0.81 0.79 3 - 0.65 0.62
With noise at level 5%
@@
With noise at level 10%
@@
KMeans experiments
Without noise
Ionosphere Banknote German Heart Balance Haberman Wholesale Indian Liver Patient Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Gini Generalized 1.18 0.4915 0.4161 4.88 0.6339 0.6363 2.48 0.6252 0.5650 1.72 0.6186 0.3495 1.84 0.8214 0.5520 2.14 0.4772 0.3409 2.36 0.4466 0.1273 1.62 0.7052 0.0759 Gini - 0.3897 0.3535 - 0.5822 0.5816 - 0.4645 0.5360 - 0.4916 0.3330 - 0.7131 0.5408 - 0.4082 0.3475 - 0.1467 0.1273 - 0.6828 0.1174 Euclidean - 0.3147 0.2882 - 0.6077 0.6115 - 0.4123 0.5130 - 0.4239 0.2672 - 0.6977 0.5296 - 0.3183 0.3213 - 0.5445 0.1227 - 0.4288 0.0155 Manhattan - 0.3586 0.3280 - 0.6083 0.6123 - 0.4123 0.5130 - 0.4113 0.2573 - 0.7204 0.5728 - 0.3247 0.3443 - 0.5992 0.0977 - 0.4288 0.0155 Minkowski 3.0 0.3200 0.2910 3.0 0.6085 0.6123 3.0 0.4123 0.5130 3.0 0.4254 0.2605 3.0 0.6475 0.4928 3.0 0.3208 0.2909 3.0 0.5315 0.1386 3.0 0.4288 0.0155 Cosine - 0.4013 0.3621 - 0.5974 0.6006 - 0.3084 0.4840 - 0.5479 0.2934 - 0.7292 0.5088 - 0.3259 0.3770 - 0.5952 0.4886 - 0.6376 0.1575 Canberra - 0.5419 0.4505 - 0.5901 0.5948 - 0.5435 0.5250 - 0.5651 0.2013 - 0.4913 0.3744 - 0.6109 0.4388 - 0.2348 0.1136 - 0.6864 0.1798 Hellinger - 0.5754 0.5021 - 0.6015 0.5868 - 0.3364 0.4740 - 0.4159 0.1985 - 0.5745 0.3760 - 0.4612 0.4304 - 0.3319 0.2273 - 0.7150 0.0414 Jensen-Shannon - 0.1289 0.3590 - 0.3085 0.5554 - 0.3095 0.5050 - 0.2548 0.0363 - 0.4687 0.1808 - 0.2361 0.0230 - 0.2621 0.1523 - 0.7150 0.0588 PearsonChi2 - 0.1289 0.3590 - 0.5043 0.5094 - 0.2789 0.4960 - 0.4802 0.2644 - 0.4578 0.3616 - 0.5301 0.4550 - 0.5397 0.4955 - 0.6623 0.1678 VicisSymmetric1 - 0.5412 0.5070 - 0.5225 0.5109 - 0.2954 0.4970 - 0.3202 0.1551 - 0.4240 0.3840 - 0.5324 0.4189 - 0.3484 0.1250 - 0.5064 0.4039 Hassanat - 0.4226 0.3821 - 0.5556 0.5517 - 0.5754 0.5750 - 0.5199 0.4033 - 0.5938 0.4464 - 0.5128 0.4769 - 0.3946 0.3250 - 0.6958 0.1400
Iris Wine Breast-Cancer Sonar QSAR Vehicle Australian Glass Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Gini Generalized 3 3.52 0.0998 0.1 3 4.10 0.1913 0.1738 2 2.70 0.9234 0.9157 2 1.22 0.6339 0.5956 3 3.12 0.3741 0.2978 3 4.52 0.2946 0.2801 2 1.66 0.7693 0.6145 6 2.54 0.1492 0.0699 Gini 3 2 0.0828 0.1 3 2 0.1172 0.1005 2 2 0.9129 0.9052 2 2 0.5985 0.5618 3 2 0.3419 0.2644 3 2 0.2596 0.2517 2 2 0.6891 0.6101 6 2 0.0584 0.0327 Euclidean 3 - 0.0757 0.0933 3 - 0.0610 0.0502 2 - 0.8802 0.8524 2 - 0.5636 0.5373 3 - 0.3520 0.2349 3 - 0.2088 0.2269 2 - 0.4873 0.5590 6 - 0.0388 0.0186 Manhattan 3 - 0.0788 0.1 3 - 0.0680 0.0613 2 - 0.8782 0.8489 2 - 0.5818 0.5518 3 - 0.3531 0.2567 3 - 0.2057 0.2234 2 - 0.4873 0.5590 6 - 0.0628 0.0279 Minkowski 3 3 0.0774 0.0933 3 3 0.0643 0.0502 2 3 0.8812 0.8542 2 3 0.5784 0.5419 3 3 0.3498 0.2311 3 3 0.2100 0.2281 2 3 0.4873 0.5590 6 3 0.0359 0.0186 Cosine 3 - 0.0188 0.02 3 - 0.1370 0.1119 2 - 0.9128 0.9051 2 - 0.5627 0.5325 3 - 0.3945 0.3055 3 - 0.1984 0.2186 2 - 0.6770 0.6725 6 - 0.0599 0.0280 Canberra 3 - 0.0386 0.0467 3 - 0.3990 0.3430 2 - 0.9237 0.9227 2 - 0.5407 0.5225 3 - 0.4449 0.3979 3 - 0.2292 0.2234 2 - 0.5208 0.5203 6 - 0.1123 0.0421 Hellinger 3 - 0.2906 0.2933 3 - 0.2646 0.2740 2 - 0.7313 0.7156 2 - 0.3424 0.4901 3 - 0.5476 0.3854 3 - 0.3701 0.3275 2 - 0.6701 0.5725 6 - 0.1579 0.1728 JensenShannon 3 - 0.4642 0.46 3 - 0.0543 0.1352 2 - 0.1272 0.3268 2 - 0.5549 0.5531 3 - 0.5241 0.5378 3 - 0.3695 0.4196 2 - 0.6830 0.6319 6 - 0.0 0.0 PearsonChi2 3 - 0.2688 0.2867 3 - 0.4505 0.4865 2 - 0.8890 0.8683 2 - 0.5973 0.5823 3 - 0.3998 0.2990 3 - 0.3644 0.3509 2 - 0.6705 0.5913 6 - 0.2025 0.1457 VicisSymmetric1 3 - 0.0667 0.0333 3 - 0.2341 0.2037 2 - 0.8656 0.8420 2 - 0.4488 0.5296 3 - 0.4766 0.2786 3 - 0.4819 0.4031 2 - 0.6502 0.5652 6 - 0.1515 0.1260 Hassanat 3 - 0.0497 0.06 3 - 0.3990 0.3598 2 - 0.9240 0.9157 2 - 0.5834 0.5470 3 - 0.3772 0.2490 3 - 0.3563 0.2931 2 - 0.7858 0.7739 6 - 0.0553 0.0279